MainPage:Nuclear:KaonDetector:PMTCharacteristics:LogBook:pag03
Jump to navigation
Jump to search
⇐ Back to the PMT Characteristics page |
⇒ Go to Summer 2012 work page |
LogBook on the CODA data acquisition, at CUA
Page (last), [6], [5], [4], 3, [2], [1].
Run Number | Date | User | PMT | High Voltage | Experiment | Description and observations | Image |
---|---|---|---|---|---|---|---|
909 (part 2) | 08/06/2012 | Laura | 9650 (BLAST) | -1.8 | Prototype Testing (Cosmic) | This is the constant DC signal that is used to track any movement in the gate for the prototype. Notice the small patch of events above the main pedestal. | Error creating thumbnail: Unable to save thumbnail to destination
|
910 | 08/06/2012 | Laura | 9650 (BLAST) | -1.8 | Prototype Testing (Cosmic) | 9cm Aerogel run in the dark-box. Using separate HV sources. Tracker still shows two patches of events above the pedestal. Notice large error on fit of signal peak. Npe~9.4 | Error creating thumbnail: Unable to save thumbnail to destination
|
910 (part 2) | 08/06/2012 | Laura | 9650 (BLAST) | -1.8 | Prototype Testing (Cosmic) | This is the constant DC signal that is used to track any movement in the gate for the prototype. Notice the small patch of events above the main pedestal. | Error creating thumbnail: Unable to save thumbnail to destination
|
911 | 08/06/2012 | Laura | 9650 (BLAST) | -1.8 | Prototype Testing (Cosmic) | 8cm Aerogel run in the dark-box. Using separate HV sources. Tracker still shows a patch of events above the pedestal. Npe~7.7 | Error creating thumbnail: Unable to save thumbnail to destination
|
911 (part 2) | 08/07/2012 | Laura | 9650 (BLAST) | -1.8 | Prototype Testing (Cosmic) | This is the constant DC signal that is used to track any movement in the gate for the prototype. Notice the small patch of events above the main pedestal. | Error creating thumbnail: Unable to save thumbnail to destination
|
912 | 08/07/2012 | Laura | 9650 (BLAST) | -1.8 | Prototype Testing (Cosmic) | 7cm Aerogel run in the dark-box. Using separate HV sources. Overnight run. Tracker: about 3 patches of signal above pedestal. Signal: mostly one signal peak (second very small patch at ~400 seems more like trailing peak). Npe~6.8 | Error creating thumbnail: Unable to save thumbnail to destination
|
912 (part 2) | 08/07/2012 | Laura | 9650 (BLAST) | -1.8 | Prototype Testing (Cosmic) | This is the constant DC signal that is used to track any movement in the gate for the prototype. Notice the small patch of events above the main pedestal. | Error creating thumbnail: Unable to save thumbnail to destination
|
913 | 08/08/2012 | Laura | 9650 (BLAST) | -1.8 | Prototype Testing (Cosmic) | 6cm Aerogel run in the dark-box. Using separate HV sources. 6 hour run. Seems to be too few data points and therefore a bad fit to the signal peak. Run 914 is the same but for a longer time. | Error creating thumbnail: Unable to save thumbnail to destination
|
913 (part 2) | 08/08/2012 | Laura | 9650 (BLAST) | -1.8 | Prototype Testing (Cosmic) | This is the constant DC signal that is used to track any movement in the gate for the prototype. Notice the small patch of events above the main pedestal. | Error creating thumbnail: Unable to save thumbnail to destination
|
914 | 08/09/2012 | Laura | 9650 (BLAST) | -1.8 | Prototype Testing (Cosmic) | 6cm Aerogel run in the dark-box. Using separate HV sources. Tracker: very small patch of signals between 400 and 600. Npe~6 | Error creating thumbnail: Unable to save thumbnail to destination
|
914 (part 2) | 08/09/2012 | Laura | 9650 (BLAST) | -1.8 | Prototype Testing (Cosmic) | This is the constant DC signal that is used to track any movement in the gate for the prototype. Notice the small patch of events above the main pedestal. | Error creating thumbnail: Unable to save thumbnail to destination
|
915 | 08/13/2012 | Laura | 9650 (BLAST) | -1.8 | Prototype Testing (Cosmic) | 5cm Aerogel run in the dark-box. Using separate HV sources. Notice small number of events across whole ADC spectrum and large overflow channel peak. | Error creating thumbnail: Unable to save thumbnail to destination
|
915 (part 2) | 08/06/2012 | Laura | 9650 (BLAST) | -1.8 | Prototype Testing (Cosmic) | This is the constant DC signal that is used to track any movement in the gate for the prototype. Notice the small patch of events above the main pedestal. | Error creating thumbnail: Unable to save thumbnail to destination
|
936 | 08/20/2012 | Marco | 60158 (HMS) | -1.810 | PMT Scanning | Scanning setup changed. Since the optical filter was broken, it was reglued to the reference PMT. LED connector was improved, reducing chances of bad connection. Doubt about the quality of the Signal cable. Scanning for test of the setup. Pulse pot high 7.5 (at +5V) and keep width 4 (at 15ns). Scanning with 10x10x1k. Reference PMT (Ref_936). | Error creating thumbnail: Unable to save thumbnail to destination
|
937 | 08/20/2012 | Marco | 60158 (HMS) | -1.810 | PMT Scanning | Running a detailed scan to check setup. Same configuration as above, just changing spacial resolution: 100x100 matrix with 1k points in each position. (Ref_937). Perhaps signal cable has a bad contact. | Error creating thumbnail: Unable to save thumbnail to destination
|
938 | 08/21/2012 | Marco | 60158 (HMS) | -1.800 | PMT Scanning | Same as run #937. Exchanging Signal cable with Trigger cable, to analyze the suspect of a bad contact in the Signal cable. Reference PMT: (Ref_938). Run stop before conclusion. | Error creating thumbnail: Unable to save thumbnail to destination
|
939 | 08/22/2012 | Marco | 60158 (HMS) | -1.800 | PMT Scanning | Same as run #938. Reference PMT: (Ref_939). PMT was too far from the LED (~5cm) and colimator was durty (silicon). | Error creating thumbnail: Unable to save thumbnail to destination
|
940 | 08/22/2012 | Marco | 60158 (HMS) | -1.800 | PMT Scanning | PMT 1cm from the colimator. Reference PMT: (Ref_940). Problems with the cable inside the box. It moved the PMT. | Error creating thumbnail: Unable to save thumbnail to destination
|
941 | 08/23/2012 | Marco | 60158 (HMS) | -1.800 | PMT Scanning | Same as run #940 (cables position fixed). Reference PMT: (Ref_941). Good run! Still need calibration to number of photoelectrons... | Error creating thumbnail: Unable to save thumbnail to destination
|
942 | 08/23/2012 | Marco | (Blast) | -1.800 | PMT Scanning | Same as run #941, changing PMT. Reference PMT: (Ref_942). Still need calibration to number of photoelectrons... Still have problems with the seput. The reference signal also show that the problem now is with the LED. I suspect that one of the cables connector is not good (the one connected to the LED, in the LED connection itself), so I tried to fix it for next run. | Error creating thumbnail: Unable to save thumbnail to destination
|
943 | 08/24/2012 | Marco | (Blast) | -1.800 | PMT Scanning | Same as run #942, trying to fix cable connected to LED. Reference PMT: (Ref_943). Scan stopped before just to analyze the conditions. Good run, setup seems to be fixed now. Still need calibration to number of photoelectrons... | Error creating thumbnail: Unable to save thumbnail to destination
|
944 | 08/24/2012 | Marco | (Blast) | -1.800 | PMT Scanning | Same as run #942. CODA crashed. | - |
950 | 08/27/2012 | Marco | (Blast) | -1.800 | PMT Scanning | Same as run #942. Reference PMT: (Ref_950). Still with problems in the scanning setup... seems to have bad connection to LED and also the scanned PMT moved? | Error creating thumbnail: Unable to save thumbnail to destination
|
951 | 08/27/2012 | Marco | (Blast) | -1.800 | PMT Scanning | Trying to fix problem on light intensity (see inomogeneity on rerefence PMT from run #950). Changed connection to LED, replacing old BNC cable by the red/green cables (soldered direct to LED, without connector anymore). Reference PMT: (Ref_951). | Error creating thumbnail: Unable to save thumbnail to destination
|
961 | 08/31/2012 | Marco | 60158 (HMS) | 0.0 | PMT rate | Running PMT with LED flashing in the center, with different rates. Rate: 100.4 Hz | Error creating thumbnail: Unable to save thumbnail to destination
|
962 | 08/31/2012 | Marco | 60158 (HMS) | -1.796 | PMT rate | Running PMT with LED flashing in the center, with different rates. Rate: 100.4 Hz | Error creating thumbnail: Unable to save thumbnail to destination
|
963 | 08/31/2012 | Marco | 60158 (HMS) | -1.796 | PMT rate | Running PMT with LED flashing in the center, with different rates. Rate: 1,024.6 Hz | Error creating thumbnail: Unable to save thumbnail to destination
|
964 | 08/31/2012 | Marco | 60158 (HMS) | -1.796 | PMT rate | Running PMT with LED flashing in the center, with different rates. Rate: 10,285.3 Hz | Error creating thumbnail: Unable to save thumbnail to destination
|
965 | 08/31/2012 | Marco | 60158 (HMS) | -1.796 | PMT rate | Running PMT with LED flashing in the center, with different rates. Rate: 108,612.1 Hz | Error creating thumbnail: Unable to save thumbnail to destination
|
966 | 08/31/2012 | Marco | 60158 (HMS) | -1.796 | PMT rate | Running PMT with LED flashing in the center, with different rates. Rate: 991,350.0 Hz | Error creating thumbnail: Unable to save thumbnail to destination
|
967 | 08/31/2012 | Marco | 60158 (HMS) | -1.796 | PMT rate | Running PMT with LED flashing in the center, with different rates. Rate: 990,980.7 Hz | Error creating thumbnail: Unable to save thumbnail to destination
|
990 | 09/04/2012 | Marco | 09577 (Blast) | -1.792 | PMT rate | Running PMT with LED flashing in the center, with different rates. Rate: 100.5 Hz | Error creating thumbnail: Unable to save thumbnail to destination
|
991 | 09/04/2012 | Marco | 09577 (Blast) | -1.792 | PMT rate | Running PMT with LED flashing in the center, with different rates. Rate: 1,026.5 Hz | Error creating thumbnail: Unable to save thumbnail to destination
|
992 | 09/04/2012 | Marco | 09577 (Blast) | -1.792 | PMT rate | Running PMT with LED flashing in the center, with different rates. Rate: 10,399.6 Hz | Error creating thumbnail: Unable to save thumbnail to destination
|
993 | 09/04/2012 | Marco | 09577 (Blast) | -1.792 | PMT rate | Running PMT with LED flashing in the center, with different rates. Rate: 108,806.1 Hz. Cutting "strange" values using reference PMT (accepting values where ref > 27). | Error creating thumbnail: Unable to save thumbnail to destination
|
994 | 09/04/2012 | Marco | 09577 (Blast) | -1.792 | PMT rate | Running PMT with LED flashing in the center, with different rates. Rate: 992,750.0 Hz. Cutting "strange" values using reference PMT (accepting values where ref > 27). | Error creating thumbnail: Unable to save thumbnail to destination
|
995 | 09/04/2012 | Marco | 09577 (Blast) | 0.0 | PMT rate | Running PMT with LED flashing in the center, with different rates. Rate: 992,502.8 Hz. Cutting "strange" values using reference PMT (accepting values where ref > 27). | Error creating thumbnail: Unable to save thumbnail to destination
|
996 | 09/04/2012 | Marco | 09577 (Blast) | 0.0 | PMT rate | Running PMT with LED flashing in the center, with different rates. Rate: 100.5 Hz. Cutting "strange" values using reference PMT (accepting values where ref > 27). | Error creating thumbnail: Unable to save thumbnail to destination
|
1011 | 09/05/2012 | Laura | 9650 (BLAST) | -1.8 | Prototype Testing (1.02) | 10cm of 1.02 aerogel in the prototype. ~24hour test. Notice low peak, was difficult to fit. | Error creating thumbnail: Unable to save thumbnail to destination
|
1012 | 09/06/2012 | Laura | 9650 (BLAST) | -1.8 | Prototype Testing (1.02) | 8cm of 1.02 aerogel in the prototype. ~24hour test. Again, notice low peak, was difficult to fit. | Error creating thumbnail: Unable to save thumbnail to destination
|
1013 | 09/07/2012 | Laura | 9650 (BLAST) | -1.8 | Prototype Testing (1.02) | 5cm of 1.02 aerogel in the prototype. ~72 test. Ran over the weekend to try to see a more distinct peak with more events; peak was still low and spread out. | Error creating thumbnail: Unable to save thumbnail to destination
|
1071 | 10/05/2012 | Marco | 9598 (BLAST) | -1.5 | Gain | As this PMT was removed from the Kaon Aerogel Detector, we are testing its gain to analyze eventual changes in its characteristics. It is important to note that this PMT was submitted to temperatures up to 150 degrees Celsius during the tentatives of its remove. LED at 1kHz. | - |
1072 | 10/05/2012 | Marco | 9598 (BLAST) | -1.6 | Gain | Same as run #1071, just changing HV. | - |
1073 | 10/05/2012 | Marco | 9598 (BLAST) | -1.8 | Gain | Same as run #1071, just changing HV. | - |
1074 | 10/05/2012 | Marco | 9598 (BLAST) | -2.0 | Gain | Same as run #1071, just changing HV. Box opened to turn reference PMT off. | - |
1075 | 10/05/2012 | Marco | 9598 (BLAST) | -1.5 | Gain | Running again from beginning of HV (see run #1071). The box was not covered with the black cover. Increased light intensity. | Error creating thumbnail: Unable to save thumbnail to destination
|
1076 | 10/05/2012 | Marco | 9598 (BLAST) | -1.6 | Gain | Same as run #1076, just changing HV.. | Error creating thumbnail: Unable to save thumbnail to destination
|
1077 | 10/05/2012 | Marco | 9598 (BLAST) | -1.8 | Gain | Same as run #1076, just changing HV.. | Error creating thumbnail: Unable to save thumbnail to destination
|
1078 | 10/05/2012 | Marco | 9598 (BLAST) | -2.0 | Gain | Same as run #1076, just changing HV.. | Error creating thumbnail: Unable to save thumbnail to destination
|
1079 | 10/05/2012 | Marco | 9598 (BLAST) | -2.2 | Gain | Same as run #1076, just changing HV.. | Error creating thumbnail: Unable to save thumbnail to destination
|
1083 | 10/05/2012 | Marco | 9598 (BLAST) | - | Gain | For attenuation factor analysis. Signal with direct pulse, not passing through all cables (approx. -400mV@10ns, similar to PMT pulse). | Error creating thumbnail: Unable to save thumbnail to destination
|
1084 | 10/05/2012 | Marco | 9598 (BLAST) | - | Gain | For attenuation factor analysis. Signal with attenuated pulse, passing through all signal cables (exactly same pulse as run #1082). | Error creating thumbnail: Unable to save thumbnail to destination
|
1085 | 10/05/2012 | Marco | 9598 (BLAST) | - | Gain | Repeating run #1083 (direct pulse). | Error creating thumbnail: Unable to save thumbnail to destination
|
1086 | 10/05/2012 | Marco | 9598 (BLAST) | - | Gain | No signal (to measure zero in the attenuation analysis). | Error creating thumbnail: Unable to save thumbnail to destination
|
- | 10/05/2012 | Marco | 9598 (BLAST) | - | Gain | Analysis of the previous measurements (from run #1075 to #1079, considering attenuation factor measured using runs #1083 to #1086. Analysis file: PMT9598_gain.xls | Error creating thumbnail: Unable to save thumbnail to destination
|
1087 | 11/07/2012 | Marco | 60158 (HMS) | -1.80 (Camac) | Prototype | 08 aerogel tiles of 1.030. Tiles cleaned with brush to improve transparency. Mylar covering inside prototype. Triggering 2in PMTs with -1.9kV. | File:1087.png |
1088 | 11/07/2012 | Marco | 60158 (HMS) | -1.80 (Camac) | Prototype | Repeating run #1087. Triggering 2in PMTs with -1.9kV. 1690 events in ~15.75 hours. | File:1088.png |
1089 | 11/07/2012 | Marco | 09650 (Blast) | -1.80 (Camac) | Prototype | Same as run #1087, changing PMT. | File:1089.png |
1090 | 11/08/2012 | Marco | 09650 (Blast) | -1.80 (Camac) | Prototype | Same as run #1089. Bottom triggering PMT stoped working at some point, but acquired data is good. | File:1090.png |
1091 | 11/13/2012 | Marco | 09650 (Blast) | -1.80 (Camac) | Prototype | Same as run #1089. Triggering PMT fixed. From 17:45 (11/13) to 14:33 (11/19) with 15399 events. | ![]() |
1092 | 11/19/2012 | Marco | 60158 (HMS) | -1.80 (Camac) | Prototype | Same as run #1091, changing PMT. From 14:50 (11/19) to 07:17 (11/26) with 12968 events. | ![]() |
1093 | 11/26/2012 | Marco | 09650 (Blast) | -1.80 (Camac) | Prototype | 08 aerogel tiles of 1.015, boxes 01...04 (01 in the bottom). Changing PMT. From 14:13 (11/26) to 17:33 (11/27) with 2833 events. Figured out that the bottom scintillator was broken (not-glued to PMT). | Error creating thumbnail: Unable to save thumbnail to destination
|
1094 | 12/03/2012 | Marco | 09650 (Blast) | -1.80 (Camac) | Prototype | Same as run #1093, scintillator fixed. From 10:18 (12/03) to 08:24 (12/05) with 5707 events. | ![]() |